Operations Report – May 18th

Crew 281 Operations Report 18-05-2023

SOL: 4

Name of person filing report: Megan Kane (Commander)

Non-nominal systems:

Lower deck shopvac filter

Electrical outlets in kitchen

Remote observatory

EVA Helmet

Notes on non-nominal systems:

No new notes.

ROVERS

Rover numbers from previous sol summaries as there was no EVA today.

Spirit rover used: No

Hours: (before EVA): 222.2

Hours: (after EVA): NA

Beginning charge: (Before EVA): NA

Ending charge: (On return from EVA, before recharging): NA

Currently charging: Yes

Opportunity rover used: No

Hours: (before EVA): 129.6

Hours: (after EVA):

Beginning charge: (Before EVA): NA

Ending charge: (On return from EVA, before recharging): NA

Currently charging: Yes

Curiosity rover used: No

Hours: (before EVA): 239.8

Hours: (after EVA): NA

Beginning charge: (Before EVA): NA

Ending charge: (On return from EVA, before recharging): NA

Currently charging: Yes

Perseverance rover used: No

Hours: (before EVA): 264.3

Hours: (after EVA): NA

Beginning charge: (Before EVA): NA

Ending charge: (On return from EVA, before recharging): NA

Currently charging: Yes

General notes on rovers: no rovers were used

Summary of Hab operations: put notes here

WATER USE: 29.4 gallons

Water (static tank): 396.4 gallons

Static tank pipe heater (on or off): off

Static tank heater (On or off): off

Toilet tank emptied: yes

Summary of internet:

nominal

Summary of suits and radios: All suits and radios are fully recharged.

Summary of GreenHab operations:

WATER USE: 1.85 gallons

Heater: Off

Supplemental light: Off

Harvest: in grams : 0

Summary of Science Dome operations:

Dual split: Heat and AC, Off

Cacao Experiment:

Pulled out and checked cacao seeds

Ham Radio Work:

Worked on software defined radios

Summary of RAM operations: (Tools used, work done)

Retrieved supplies for projects in Hab and Greenhab

Summary of any observatory issues: NA

Summary of health and safety issues:

No new issues to report.

Questions, concerns and requests to Mission Support:

Request Mission Support assistance for the remainder of the mission related to the Ham Radio antenna being installed tomorrow in case of emergency or damage.

Sol Summary – May 17th

Crew 281 Sol Summary Report 17 May 2023
Sol: 3
Summary Title: Geotechnical Surveys and Bread
Author’s name: Ritupriya Patil (XO) and Megan Kane (Commander)
Mission Status: Nominal
Sol Activity Summary:
EVA 3: Ana Pires (Crew Scientist)(EVA Lead) and KC Shasteen (Crew Engineer)(EVA Comm). Hab Comm: Ritupriya Patil (XO)
EVA prep began at 7 am when we started prepping the suits, radio comms, helmets and air circulation. KC Shasteen (Crew Engineer) made the pre-EVA checks with radios and spacesuits. After a 15 minute pre-breath followed a 5 minute depressurization in the air lock, leaving the airlock at 7:35 am. Waterchecks were scheduled after each station in the Geotechnical survey. Mid EVA one piece of equipment required repair, it was completed successfully. All objectives were completed. The Geotechnical survey took less time than initially expected and the team returned early.
In the Greenhab an airflow diverter was tested. With this change there should be no future concerns during this mission about excessive airflow on the plants in front of the fan when it is in use.
In the science dome, 2 research projects were worked on.
Lunch was freshly baked bread with butter and a delicious sauteed veggie medley. Ritupriya (XO) held a yoga stretch session to destress and work on the post EVA aches and soreness.
Dinner was left overs from all the yummy meals in the past couple days.
Look Ahead Plan:
Tomorrow we plan to have a restful day. We will do hab operations, maintenance, and daily research activities and set aside time for recreational activities. There will be no EVA tomorrow.
Anomalies in work: None
Weather: Windy, 39 Km/h gusts
Crew Physical Status: No changes
EVA: Completed 1 today
Reports to be filed:
● Greenhab Report
● EVA Report
● Ops Report
● Journalist Report
● PHOTOS!
● Photo of the Day
Support Requested:
● None

GreenHab Report – May 17th

Crew 281 GreenHab Report 17 May 2023
GreenHab Officer: Megan Kane
Environmental control: fan with door open
Average temperatures: 90.38
Hours of supplemental light: None
Daily water usage for crops: 1.72 gallons
Daily water usage for research and/or other purposes: NA
Water in Blue Tank 174.19 gallons
Time(s) of watering for crops:
0800: refilled 8 water bottle (4 L) watered other plants as needed 1.5 L
1505: refilled 1 water bottle (500 mL) watered other plants as needed 500 mL
Changes to crops:
Average temperature was the highest it has been since we started. The highest tomato leaves dropped a little during the hottest part of the day but there is no sign of wilting. The plants are in good condition.
Narrative:
In the morning the greenhab was checked and watered as required. For plants without a passive watering system, a dowel rod was used to perforate the soil crust in 1-3 places per planter to allow the water to get down to the roots more efficiently and reduce evaporation.
The morning was a bit hazy and the temperature during the initial test of the airflow diverter went well. However once the haze dissipated the interior quickly rose in temperature. Opening the door after the test had a slight effect. The Crew Engineer and Greenhab officer looked at the evaporative cooling system to see if it could be utilized as we have entered the time of year and weather conditions where it would be effective. It was determined that we would need to reach out to mission control and that setup would likely require an EVA.
This afternoon ingredients for making a marinara sauce were harvested.
There is concern about the lack of humidity in the greenhab and the increasing temperatures.
Harvest: (include which crop and mass in grams)
Tomatoes 514
Basil 1
Sage 1
Chives 7
Oregano 1
Thyme 1
Support/supplies needed:
We request additional information on evaporative cooler including:
· Make and model
· Date last used
· Status when last used (operational or broken)
· Method of connecting the system to a water supply and which one.

EVA Report – May 17th

Crew 281 EVA Report 17-05-2023

EVA # 4

Authors: Ritupriya Patil (Executive Officer), Ana Pires (Crew Scientist)

Purpose of EVA:
Conduct Geotechnical Survey of the area and geomechanical assessment
Schmidt Hammer and Equotip measurements along scanlines
Geological and geotechnical assessment of the scanline using the created datasheet by the Crew Scientist
Acquire images and collect loose rock/soil
Take pictures at Marble Ritual, on the way back, if time and energy allow

Start time: 0735 hrs

End time: 0944 hrs

Narrative:
Crew of started pre-EVA prep at 7 am. KC Shashteen (Crew Engineer) did radio and suit checks. The crew donned the suits and were ready for 15 mins of pre-breathing at 0715 hrs. The crew left the hab at 0735 hrs to begin the EVA on Opportunity rover, which was driven to Pooh’s corner. The crew arrived at Pooh’s corner at 0748 hrs and found a location for rock and soil sample collection about 125 m north of where the rover was parked. They began hardness sampling with a scanline of 200 m, and set up fibreglass tape, Schimdt hammer and equotip equipment. The crew took water breaks to keep themselves hydrated throughout the EVA. After a 5 station long scanline they completed the measurements and returned to the vehicle at 0913 hrs. On their way back to the hab, the crew stopped at Marble Ritual for some pictures of the geological features.
Overall, the crew had a successful EVA with new learnings and they re-entered the hab at 0944 hrs after 5 mins of pressurisation time in the air lock.

Overview of EVA

Collect rock and soil samples for hardness analysis
Visual observation on the area at Marble Ritual

Issues Experienced During EVA

The crew had only one pen to record the scanline measurements. This became problematic when the ball point was badly damaged since it may have hit a rock and it temporarily stopped working. They were able to make it work but we should have a contingency plan.
The equipment came unscrewed and due to the gloves, the crew lost the sensation that the equipment fell off. They were able to locate the part and put it back together to continue with measurements.
Tourists showed up at the EVA location again but there was no interaction from them.

Outcomes

The crew had some confusion on first stop for the EVA, there should have been a pre-brief in the morning that did not happen.
The system works well but we need to understand how to integrate the geotechnical equipment with the spacesuit and gloves for better handling in future space activities.
Collected hardness sampling measurements for 5 station long scanlines.
Collected 18 rocks of 1120 grams and 1 soil sample of 95 grams at Pooh’s corner

Destination: a) Pooh’s Corner and walking no more than a radius of 500m; b) On the way back, stop in Marble Ritual and walk no more than 300m.
Coordinates a) (use UTM WSG 84): 0519025 E, 4251183 N
Coordinates b) (use UTM WSG 84): 0518687 E, 4250789 N
EVA Participants 2: KC (Crew Engineer), Ana Pires (Crew Scientist and EVA Leader)
Road(s) and routes per MDRS Map: Drive on Cow Dung Road and walk east until Pooh’s Corner. On the way back, a quick stop in Marble Ritual to take extra photos.
Mode of travel: Opportunity

Journalist Report – May 17th

Crew 281 – Journalist Report, 17 May 2023

Report by: Rachel Jones

Sol 3, Today was our first "longer" EVA mission. The whole team worked to prepare KC and Ana. They were going to Pooh’s Corner and Marble Ritual to perform scan lines while testing new geotechnical equipment. Once the EVA team returned everyone helped them doff their gear.

For lunch Megan made fresh bread. We ate bread dipped/spread/stuffed with a sauteed vegetable medley containing tomatoes harvested from the Green Hab combined with freeze dried vegetables and rice vinegar. It was astonishingly delicious. We’ve voted that the rice vinegar KC brought was a welcome edition to the MDRS kitchen.

The afternoon sent us all to our various projects. Megan checked on her chocolate seeds. Ana analyzed her collected data. Ritu started on our reports and planned for the next EVA. KC handled various Hab maintenance. We had strong winds in the afternoon and KC fortified the tunnel system. I (not having cooked) challenged myself to use less water doing dishes. MDRS recommends a sink of soapy water and a sink of rinse water. Then I did general Hab cleaning on the upper floor, before taking an inventory of MDRS’s amateur radio station. Truly, it’s science, maintenance, and cleaning everyday.

Pre-dinner we did a group yoga session. Dinner tonight is leftovers. I’m hoping to eat some of Sol 2’s quesadillas.

Operations Report – May 17th

Crew 281 Operations Report 17-05-2023

SOL: 3

Name of person filing report: Megan Kane

Non-nominal systems:

· Lower deck shopvac

· Electrical outlets in kitchen

· Remote observatory

· EVA Helmet

Notes on non-nominal systems:

· No new notes.

ROVERS

Spirit rover used: No

Hours: (before EVA): NA

Hours: (after EVA): NA

Beginning charge: (Before EVA): NA

Ending charge: (On return from EVA, before recharging): NA

Currently charging: Yes

Opportunity rover used: Yes

Hours: (before EVA): 129.4

Hours: (after EVA): 129.6

Beginning charge: (Before EVA): 100

Ending charge: (On return from EVA, before recharging): 94

Currently charging: Yes

Curiosity rover used: No

Hours: (before EVA): NA

Hours: (after EVA): NA

Beginning charge: (Before EVA): NA

Ending charge: (On return from EVA, before recharging): NA

Currently charging: Yes

Perseverance rover used: No

Hours: (before EVA): NA

Hours: (after EVA): NA

Beginning charge: (Before EVA): NA

Ending charge: (On return from EVA, before recharging): NA

Currently charging: Yes

General notes on rovers: Opportunity was used today

Summary of Hab operations: put notes here

WATER USE: 32.7 gallons

Water (static tank): 425.8 gallons

Static tank pipe heater (on or off): off

Static tank heater (On or off): off

Toilet tank emptied: no or yes: no

Summary of internet:

· Open internet hours now from 6 to 10 pm

Summary of suits and radios: All suits and radios are fully recharged.

Summary of GreenHab operations:

WATER USE: 1.72 gallons

Heater: Off

Supplemental light: Off

Harvest:

Tomatoes 514

Basil 1

Sage 1

Chives 7

Oregano 1

Thyme 1

Summary of Science Dome operations:

Dual split: Heat and AC, Off

Cacao Experiment:

Cacao seeds in trays. Temperature and Humidity have remained in the desired range throughout the day. Seeds and trays were cleaned and set out again. Seeds that are not sprouting were removed.

Geotechnical Experiment Activities:

· Conducted tests geotechnical equipment

· Weighed samples from EVA 4

Summary of RAM operations: (Tools used, work done)

No new work in the RAM.

Summary of any observatory issues: NA

Summary of health and safety issues:

· No new issues to report.

Questions, concerns and requests to Mission Support:

· Crew would like to review procedures for in-sim contingencies, for example suit failure or a rover needs to be towed. If there are examples of lessons learned from previous crews it would be helpful to receive copies.

Sol Summary – May 16th

Crew 281 Sol Summary Report 16-05-2023
Sol: 2
Summary Title: Around the Hab and Experiments
Author’s name: Ritupriya Patil (XO) and Megan Kane (Commander)
Mission Status: Nominal
Sol Activity Summary:
Before the EVA prep began Megan Kane (Commander) checked the greenhab.
EVA 3: Ritupriya Patil (XO, EVA Lead), Ana Pires (Crew Scientist)(EVA Comm), Rachel Jones (HSO). Hab Comm: Megan Kane
EVA prep began at 7:45 am when we started prepping our suits, radio comms, helmets and air circulation. KC Shasteen (Crew Engineer) made the pre-EVA checks with radios and spacesuit. We donned the EVA team’s suits, conducted a 15 minute pre-breath followed by a 5 minute depressurization in the air lock, leaving the airlock at 8:35 am. Waterchecks were scheduled for every 45 minutes. Mid EVA one suit required adjustments. All objectives were completed. Due to the suit requiring further adjustment the EVA crew returned to the hab early at 0956. After the EVA crew returned, KC and Megan assisted with adjusting the suit for future EVAs and, hopefully, solving the issue with the fit. Additionally during the EVA there was a communications issue that was successfully resolved.
In the Greenhab an airflow diverter was added to the fan to improve the conditions for the plants.
In the science dome, 2 research projects were worked on.

Look Ahead Plan:
Tomorrow we plan to conduct the first Geotechnical survey for the Crew Scientists research.
Additional experiments are in the works and existing experiments will be continued.
Anomalies in work: None
Weather: Fine
Crew Physical Status:
· 1 crew in suboptimal condition, monitoring
EVA: Completed 1 today
Reports to be filed:
● Greenhab Report
● EVA Reports EVA 3
● Ops Report
● Commander’s Report
Support Requested:
● Issue with power system was observed and reported to Mission Support.
● Trash pickup requested for this evening.

Journalist Report – May 16th

Crew 281 Sol 2 Journalist Report 16 May 2023

Setting the Scene for Science!

By: Rachel Jones

Sol 2: 16 May 2023

Sol 2 on Mars. Today started with a multiple mission EVA. Ritupriya, Ana, and myself conducted a close range EVA around the areas immediately surrounding the habitat. Ritu conducted drone surveillance of the area. Ana looked for a suitable area to deploy her geotechnical robotic device in a later mission. My goal was to look at the area where the amateur radio antenna was previously deployed. Then a final team goal was to pick up debris surrounding the Hab. We collected 38 grams of debris! Overall it was a mission success.

Ritu was in charge of meals and hab support for Sol 2. Afternoon consisted of a wonderful Ramen lunch, followed by multiple report writing. I had cleaning duty and tried to ensure our EVA ready room remained as clean as possible. Utah dirt is not unlike Georgia clay. (Growing up on Flat Creek Ranch has prepared me well.) Megan is growing chocolate in the grow tent. Her research goal focuses on lurxary crops for space food production. (People will definitely want chocolate and Coffee on Mars.) Ana gathered all her materials for a long EVA tomorrow. KC designed and built an experimental air flow diverter for the Green Hab. Previously, Megan had rigged an experimental Terra Cotta watering system in the Green Hab. It’s taking a lot of effort from Megan and KC to keep the plants alive in the extreme Utah desert heat.

The night ended with a chickpea curry and rice dinner. (So good) Lots of discussion on future research and collaborations happened over dinner. We also thought of our team morale Dungeons and Dragons (D&D) campaign which should start soon. Finally, the evening centered on report writing. There are over 6-8 daily reports that have to be submitted, revised, then re-submitted.

Another successful day on Mars!